OpenStudy (anonymous):

logx+log(3x-13)=1 I got the answer of 10/3 can someone see if that is right??

OpenStudy (anonymous):

no it is wrong.

OpenStudy (anonymous):

Ok, thank you.

OpenStudy (anonymous):

log(3x2-13x) = 1 3x2-13x = 10

OpenStudy (anonymous):

I see what I did wrong. I put 26 instead of 13..

OpenStudy (anonymous):

Wait.. no I didnt. Then I move the 10 over right?

OpenStudy (anonymous):

giving me: 3x2-13-10=0 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Then I did the quadratic... maybe I did that wrong

OpenStudy (anonymous):

3x2-15x+2x-10

OpenStudy (anonymous):

where did you get the middle? -15x+2x?

OpenStudy (anonymous):

-13x can be represented as -15x +2x.

OpenStudy (anonymous):

why would I want to put that into my equation?

OpenStudy (anonymous):

when I have 3x2-13x-10. Would you do the quadratic then?

OpenStudy (anonymous):

is the answer 5

OpenStudy (anonymous):

it is a method to factor the quadratic equation. 3x2-13x-10 = 0 3x2-15x+2x-10 = 0 3x(x-5)+2(x-5)= 0 ,factor out x-5, (x-5)(3x+2)=0. so, x= 5 and x = -3/2

OpenStudy (anonymous):

got it!!! Thank you. I didnt notice it factored.

OpenStudy (anonymous):

you should always see if a quadratic equation can be factored, it is much easier. If that fails, do the formula to solve

OpenStudy (anonymous):

I usually do, my teacher just never gives us the ones that factor...lol.Thanks

OpenStudy (anonymous):

you are welcome